OpenStudy (anonymous):

find the limit of the following questions lim x^2/(x^2)-9 x->3 lim 4-x/(x^2)-16 x->-4 I got "does not exist" for both questions.. but I don't think it's correct can anyone help please?!

OpenStudy (anonymous):

is the first one this: \(\large\lim_{x \to 3} \frac{x^2}{x^2-9}\) ????

OpenStudy (anonymous):

if yes, then you're correct. but the second one does have a limit...

OpenStudy (anonymous):

yea that's the first one but i can't factorize the second one

OpenStudy (anonymous):

i stop over here:S \[\frac{ 4-x }{ (x-4)(x+4) }\]

OpenStudy (anonymous):

\(\large \frac{ 4-x }{ (x-4)(x+4) }=\frac{ -1(x-4) }{ (x-4)(x+4) } =\frac{ -1\cancel{(x-4)} }{\cancel{ (x-4)}(x+4) } \)

OpenStudy (anonymous):

thanks. which means i'll probably get a zero for today's test lol